OpenStudy (anonymous):

How would I further simplify the following? \(\ \large (1)(r^2+1)^{-\frac{1}{2}}+r(-\frac{1}{2}(r^2+1))^{-\frac{3}{2}}\)

OpenStudy (anonymous):

get rid of the 1 for a start

OpenStudy (anonymous):

Yup, just did that

OpenStudy (anonymous):

i bet this is some sort of derivative from a produce rule maybe

OpenStudy (anonymous):

and you took literally the power rule maybe it would be best to write as \[\frac{1}{\sqrt{r^2+1}}+\frac{r}{2\sqrt{r+1}^3}\]

OpenStudy (anonymous):

now if you want to add these up, multiply the first fraction top and bottom by \[2(r^2+1)\]

OpenStudy (anonymous):

oh it was a subtraction, no matter works the same

OpenStudy (anonymous):

@satellite73 Yup! It's from a product rule! Nice!